# Impact of Weak Lensing Mass Mapping Algorithms on Cosmology Inference

### v1

@@ 3.2 Higher-Order Statistics
In this section, we describe the HOS in our pipeline: peak counts and wavelet peak counts.
Weak lensing peaks are defined as local maxima of the convergence field; in particular, we define peaks as pixels with a value larger than that of their eight neighbours in the image. These peaks trace projected overdensities in the matter distribution, and thus often indicate the presence of massive cosmic structures, although the precise relation between peaks and halos is complex due to projection effects and potential noise-induced false detections (Sabyr et al., 2022).
Peak count analysis is performed by calculating the peak function, which represents the number of peaks in a convergence map or signal-to-noise map as a function of the peak height (the respective pixel value). In this work, we count the peaks in signal-to-noise maps, obtained by dividing the noisy convergence maps $\kappa$, convolved with a filter $\mathcal{W}_{\theta}$, by its corresponding rms map, $\sigma_{\textrm{n}}^{\mathrm{filt}}$:
@@
Here, $\ast$ denotes the convolution operator, and ${\sigma_{\rm n}^{\mathrm{filt}}}$ is derived from the rms map $r$ of $\kappa$ by ${\sigma_{\rm n}^{\mathrm{filt}}}=\sqrt{\mathcal{W}_{\theta}^{2}*r^{2}}$.
@@
The mono-scale approach consists of simply convolving the convergence maps with a single Gaussian kernel of size $\theta$. The kernel size is chosen as a compromise between the need to reduce the noise and the need to preserve the signal. The multi-scale peak count analysis is performed by employing a wavelet decomposition to the map, namely the isotropic undecimated wavelet transform, commonly referred to as the starlet transform (Starck et al., 2007). This transform enables one to extract the information encoded in different spatial scales quickly and simultaneously by decomposing the convergence maps into a set of wavelet bands of different scales. Specifically, the starlet filter decomposes the original $N\times N$ convergence map $\kappa$ into a sum of images with the same number of pixels at different resolution scales $j$:
@@
where $c_{J}$ is a coarse, very smoothed, version of the original map, and $w_{j}$ are the wavelet bands representing the details of the original map at scale $2^{j}$ pixels, $j_{\rm{max}}$ is the maximum scale considered, and $J=j_{\rm{max}}+1$. An important advantage of this analysis over other multi-scale approaches is that each wavelet band covers a different frequency range, resulting in a nearly diagonal peak count covariance matrix (Lin & Kilbinger, 2018; Ajani, 2021), which does not occur, for example, in the application of a multi-scale Gaussian analysis on the convergence map. In this starlet multi-scale approach, the standard deviation of noise $\sigma_{\rm n}^{\mathrm{filt}}$ must be assessed separately for each wavelet band. It has been shown (Starck et al., 2015) that for starlet-decomposed images, the standard deviation of the noise at scale $j$ can be estimated from the standard deviation of the noise of the original image $\sigma_{\textrm{I}}$ and some coefficients $\sigma^{e}_{j}$ that represent the behaviour of the noise in wavelet space, as $\sigma_{j}=\sigma^{e}_{j}\sigma_{\textrm{I}}$. The coefficients $\sigma^{e}_{j}$ are equal to the standard deviation at scale $j$ of the starlet transform of a normal distribution with a standard deviation of one.
@@ 4.1 Shear Maps
The output of the simulations described in Sect. 2 is a set of galaxy catalogues, each containing the positions, redshifts, the simulated shear components of the galaxies, and the simulated convergence field. We first concatenate the four redshift bins into one, so that each catalogue contains galaxies of all redshifts. We then project the galaxy right ascension and declination to Cartesian coordinates $(x,y)$ onto a grid with a pixel size of 1 arcmin. This results in $600\times 600$ pixel maps. To create the mask, we construct a galaxy number density map by binning the galaxy positions into the pixel grid, ($\textbf{M}=0$ where there are no galaxies, $\textbf{M}=1$ where there are galaxies).
@@
Next, we bin the simulated shear components in the pixel grid to create the shear maps, and generate noise maps for both shear components using
$\sigma_{\textrm{n}}^{2}=\frac{\langle\sigma_{\rm int}^{2}\rangle}{\sqrt{2\,n_{\rm gal}}},$ (5)
@@
where $\langle\sigma_{\rm int}^{2}\rangle=0.44$ is the average global intrinsic ellipticity dispersion of the source galaxies, and $n_{\rm gal}$ is the number of galaxies in the pixel. Here, we do not assume a constant galaxy number density, but rather use the actual number of galaxies in each pixel, to create a more realistic noise map. This noise map is added to the shear maps to create the mock observed shear maps.
@@
Some studies (e.g., Marques et al., 2024; Cheng et al., 2024), choose to apply some smoothing to the shear maps before the mass mapping inversion. Our approach primarily involves filtering the reconstructed mass maps with various filters before calculating the HOS (see Section 3.2). However, we also incorporate a preprocessing step where the shear maps are smoothed using a Gaussian kernel with a small standard deviation of $0.4$ arcmin. This additional shear smoothing aims to reduce edge effects in the mass maps, particularly in the KS method, caused by pixels with no shear information. As shown in Section 4.2, this smoothing slightly improves the reconstruction accuracy of certain mass maps. Since we subsequently convolve the reconstructed mass maps with larger filters, this initial smoothing is not expected to bias our results. Ultimately, we conducted our analysis with both smoothed and unsmoothed shear maps but we observed no significant difference in the resulting cosmological parameter estimation. Consequently, we report only the results obtained with shear smoothing in this paper.
4.2 Mass mapping
In order to create mass maps, we apply the KS, iKS, and MCALens methods to the noisy shear maps. For the iKS and MCALens methods, we use 50 iterations. We set the detection threshold for the peaks in MCALens to $5\sigma$.
@@ 4.2 Mass mapping
In Figure 1, we show the mass maps obtained using the KS, iKS, and MCALens methods from a single tile of the simulated footprint. The KS and iKS maps have been smoothed with a Gaussian kernel with a standard deviation of $2$ arcmin, to render the structures more visible. Without smoothing, the KS and iKS maps are dominated by noise. The MCALens maps, however, are already noise-suppressed by the mass-mapping algorithm, so they do not require additional smoothing for visualization purposes.
@@
The iKS map uses inpainting to fill in the missing data in the convergence field (see Sect. 3.1), effectively reducing the edge effects visible in the KS map. For reference, large masked regions are outlined in white in the plots of the reconstructed mass maps. Additionally, we see that the MCALens method removes the noise, while keeping the stronger peaks in the convergence field, as expected from the method.
@@
where $i$ is the pixel index, $N_{\rm mask}$ is the total number of unmasked pixels (i.e., where $\mathbf{M}_{i}=1$), and $\mathbf{M}$ is the binary mask. As shown in the table, applying Gaussian smoothing to the shear maps reduces the RMSE for the KS and iKS methods, indicating improved reconstruction accuracy. However, shear smoothing has minimal effect on the MCALens method, which consistently outperforms the others by achieving the lowest RMSE.
@@
KS (no $\gamma$ smoothing) $1.1\times 10^{-2}$
iKS (no $\gamma$ smoothing) $1.1\times 10^{-2}$
KS (with $\gamma$ smoothing) $1.0\times 10^{-2}$
iKS (with $\gamma$ smoothing) $1.0\times 10^{-2}$
MCALens (no $\gamma$ smoothing) $9.8\times 10^{-3}$
MCALens (with $\gamma$ smoothing) $9.8\times 10^{-3}$
Table 1: Root mean square error (RMSE) comparison between the true convergence map and the reconstructed mass maps obtained using the KS, iKS, and MCALens methods. The RMSE is evaluated for mass maps derived from shear maps that are either smoothed with a Gaussian kernel of standard deviation $0.4$ arcmin or left unsmoothed. The values represent the average RMSE across all realizations of the 25 cosmologies. To ensure a fair comparison, each mass map is further smoothed with a Gaussian kernel optimized to minimize its respective RMSE.
Figure 1: Mass maps obtained using different reconstruction techniques, from a specific tile of the simulated footprint. The upper left panel displays the true convergence map derived directly from the simulated $\kappa$ field. The upper right and lower left panels show the mass maps generated by the Kaiser-Squires (KS) and iterative Kaiser-Squires (iKS) methods, respectively, both smoothed with a Gaussian kernel (standard deviation of 2 arcmin) for visualization purposes. The lower right panel illustrates the MCALens mass map. The iKS method implements inpainting of the missing data, and the large masked regions of the patch are outlined in white.
@@ 4.3 Statistics calculation
Once the mass maps are generated, we calculated our HOS on $S/N$ maps, derived from the mass maps, as described in Section 3.2. We calculate the peak counts and wavelet peak counts only in regions where $\textbf{M}=1$, for both inpainted and non-inpainted maps, as the inpainted parts do not contain real signal. For both analyses, the $S/N$ maps are constructed using Eq. (3), and peaks are counted and binned into 20 linearly spaced bins over the range $-2<S/N<6$, following Ajani et al. (2020).
@@
For the mono-scale peak counts, we use as the filter $\mathcal{W}$ in Eq. (3) a Gaussian kernel. We tested various kernel sizes and selected the size that yielded the best cosmological constraints for each method. The optimal kernel sizes were $\theta_{\rm ker}=2$ arcmin for the KS and iKS methods and $\theta_{\rm ker}=1$ arcmin for the MCALens method. We then count the peaks in the resulting $S/N$ maps. Figure 2 displays the histograms of the peak counts for the KS, iKS, and MCALens methods, for the 25 simulated cosmologies. The histograms show the sensitivity of peak counts to cosmology as well as the difference between the mass-mapping methods.
@@
Figure 2: Histograms of the peak counts as a function of the signal-to-noise ratio. The three panels from left to right correspond to the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods, for the 25 simulated cosmologies. The histograms are colour-coded by their $S_{8}$ value.
@@
For the wavelet peak counts, we first decomposed the convergence maps into wavelet scales using the starlet transform, as described in Sect. 3.2. We used five wavelet scales, which correspond to resolutions $[2,4,8,16,32]$ arcmin, along with the coarse map. We then calculate the $S/N$ maps for each wavelet band separately, according to Eq. (3) and count the peaks in each $S/N$ map. Figure 3 shows the histograms of the wavelet peak counts for the KS, iKS, and MCALens methods, for the 25 simulated cosmologies. As expected, scales corresponding to smaller wavelet filters contain more peaks, as they capture smaller, more abundant structures in the convergence field. Notably, for most scales, MCALens results in a narrower distribution of peak counts, which is a result of the noise suppression by the method.
@@
Figure 3: Histograms of the wavelet peak counts as a function of the signal-to-noise ratio for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods, for the 25 simulated cosmologies. The four rows correspond to the three wavelet scales, $[8,16,32]$ arcmin and the coarse map, and the columns correspond to the three mass mapping methods. The histograms are colour coded by their $S_{8}$ value.
@@ 4.4 Covariance matrices
We calculate the covariance matrices for the HOS, using the Covariance Training Set described in Section 4.4. To avoid unphysical large-scale mode-coupling that arises from replicating the same light-cone across tiles, we randomize the selection of light-cones during the survey construction, following Harnois-Déraps et al. (2021). For each of the 124 full survey realizations, we draw the 19 tiles from different, randomly selected light-cones. We apply this process to all 10 realizations and average them to create our final covariance matrix. The covariance matrix elements are calculated as:
@@
where $N=1240$ is the number of realizations in the Covariance Training Set, $x_{i}^{r}$ is the value of the statistic in the $i$-th bin of the $r$-th realization, and $\bar{x}_{i}$ is the mean of the $i$-th bin over all realizations:
@@
We rescale the covariance matrices to account for the difference between the area of a single tile and the total survey area. Each tile covers $10\times 10$ deg2, and the survey footprint consists of 19 such tiles. To adjust for this, we divide each element of the covariance matrix by the ratio of the single-tile area to the total survey area. Moreover, to account for the finite number of bins and realizations, we use the de-biased estimator for an inverse Wishart distribution (Siskind, 1972; Hartlap et al., 2007) for the inverse of the covariance matrix:
@@
where $n_{\rm bins}$ is the number of bins in the histogram of the HOS under consideration, and $\bm{\mathrm{C}}_{*}^{-1}$ is the inverse of the respective covariance matrix calculated according to Eq. (7).
In Fig. 4, we show the correlation matrices for the single-scale peak counts, for the different mass mapping methods. Figure 5 shows the correlation matrices for the wavelet peak counts.
Figure 4: Correlation matrices for the single-scale peak counts, for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods.
Figure 5: Correlation matrices for the wavelet peak counts, for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods. The white dashed lines separate the different wavelet scales, which are positioned in increasing order of resolution from left to right, i.e. $[2^{\prime},4^{\prime},8^{\prime},16^{\prime},32^{\prime},\textrm{coarse}]$.

### v2

@@ 3.2 Higher-order statistics
In this section, we describe the HOS in our pipeline: peak counts and wavelet peak counts. Weak-lensing peaks are defined as local maxima of the convergence field; in particular, we defined peaks as pixels with a value larger than that of their eight neighbours in the image. These peaks trace projected overdensities in the matter distribution, and thus often indicate the presence of massive cosmic structures, although the precise relation between peaks and halos is complex due to projection effects and potential noise-induced false detections (Sabyr et al. 2022).
Peak count analysis is performed by calculating the peak function, which represents the number of peaks in a convergence map or signal-to-noise map as a function of the peak height (the respective pixel value). In this work we counted the peaks in signal-to-noise maps, obtained by dividing the noisy convergence maps $\kappa$, convolved with a filter $\mathcal{W}_{\theta}$, by its corresponding rms map, $\sigma_{\textrm{n}}^{\mathrm{filt}}$:
@@
Here, $\ast$ denotes the convolution operator, and ${\sigma_{\rm n}^{\mathrm{filt}}}$ is derived from the rms map $r$ of the shear field by ${\sigma_{\rm n}^{\mathrm{filt}}}=\sqrt{\mathcal{W}_{\theta}^{2}*r^{2}}$.
@@
The mono-scale approach consists of simply convolving the convergence maps with a single Gaussian kernel of size $\theta$. The kernel size is chosen as a compromise between the need to reduce the noise and the need to preserve the signal. The multi-scale peak count analysis is performed by employing a wavelet decomposition to the map, namely the isotropic un-decimated wavelet transform, commonly referred to as the starlet transform (Starck et al. 2007). This transform enables one to extract the information encoded in different spatial scales quickly and simultaneously by decomposing the convergence maps into a set of wavelet bands of different scales. Specifically, the starlet filter decomposes the original $N\times N$ convergence map $\kappa$ into a sum of images with the same number of pixels at different resolution scales $j$:
@@
where $c_{J}$ is a coarse, highly smoothed version of the original map, and $w_{j}$ are the wavelet bands representing the details of the original map at scale $2^{j}$ pixels, $j_{\rm{max}}$ is the maximum scale considered, and $J=j_{\rm{max}}+1$. An important advantage of this analysis over other multi-scale approaches is that each wavelet band covers a different frequency range, resulting in a nearly diagonal peak count covariance matrix (Lin & Kilbinger 2018; Ajani 2021), which does not occur, for example, in the application of a multi-scale Gaussian analysis on the convergence map. In this starlet multi-scale approach, the standard deviation of noise $\sigma_{\rm n}^{\mathrm{filt}}$ must be assessed separately for each wavelet band. In our analysis, we assumed that the noise in the convergence maps is non-stationary (i.e. it depends on the number of galaxies per pixel) but Gaussian, as described in Sect. 4.1. Thanks to the linearity of the starlet transform, this pixel-level noise propagates directly to the wavelet domain, allowing us to derive a spatially dependent noise level for each wavelet coefficient at every scale and position. It has been shown (Starck et al. 2015) that for starlet-decomposed images, the standard deviation of the noise at scale $j$ can be estimated from the standard deviation of the noise of the original image $\sigma_{\textrm{I}}$ and some coefficients $\sigma^{e}_{j}$ that represent the behaviour of the noise in wavelet space, as $\sigma_{j}=\sigma^{e}_{j}\sigma_{\textrm{I}}$. The coefficients $\sigma^{e}_{j}$ are equal to the standard deviation at scale $j$ of the starlet transform of a normal distribution with a standard deviation of one.
@@ 4.1 Shear maps
The output of the simulations described in Sect. 2 is a set of galaxy catalogues, each containing the positions, redshifts, simulated shear components, and simulated convergence of the galaxies. We first concatenated the four redshift bins into one, so that each catalogue contained galaxies of all redshifts. We then projected the galaxy right ascension and declination to Cartesian coordinates $(x,y)$ onto a grid with a pixel size of 1 arcmin. This resulted in $600\times 600$ pixel maps. To create the mask, we constructed a galaxy number density map by binning the galaxy positions into the pixel grid, ($\textbf{M}=0$ where there are no galaxies, $\textbf{M}=1$ where there are galaxies).
@@
Next, we binned the simulated shear components in the pixel grid to create the shear maps and generated a noise map for each shear component using
$\sigma_{\textrm{n}}^{2}=\frac{\langle\sigma_{\rm int}^{2}\rangle}{{2\,n_{\rm gal}}},$ (5)
@@
where $\langle\sigma_{\rm int}^{2}\rangle=0.44$ is the average global intrinsic ellipticity dispersion of the source galaxies, and $n_{\rm gal}$ is the number of galaxies in the pixel. Here, we did not assume a constant galaxy number density but instead used the actual number of galaxies in each pixel, to create a more realistic noise map. This noise map was added to the shear maps to create the mock observed shear maps.
@@
Some studies (e.g. Marques et al. 2024; Cheng et al. 2025) chose to apply some smoothing to the shear maps before the mass-mapping inversion. Our approach primarily involves filtering the reconstructed mass maps with various filters before calculating the HOS (see Section 3.2). However, we also incorporated a preprocessing step where the shear maps are smoothed using a Gaussian kernel with a small standard deviation of $0.4$ arcmin. This additional shear smoothing aims to reduce edge effects in the mass maps, particularly in the KS method, caused by pixels with no shear information. As shown in Section 4.2, this smoothing slightly improves the reconstruction accuracy of certain mass maps. Since we subsequently convolved the reconstructed mass maps with larger filters, this initial smoothing is not expected to bias our results. Ultimately, we conducted our analysis with both smoothed and unsmoothed shear maps but find no significant difference in the resulting cosmological parameter estimation. Consequently, we report only the results obtained with shear smoothing in this paper.
4.2 Mass-mapping
To create mass maps, we applied the KS, iKS, and MCALens methods to the noisy shear maps. For the iKS and MCALens methods, which include an internal iterative process, we set the number of iterations to 50, since this provides a good compromise between achieving sufficient convergence and managing computational time. We also set the internal parameter of the peak detection threshold in MCALens to $5\sigma$.
@@ 4.2 Mass-mapping
In Figure 1, we show the mass maps obtained using the KS, iKS, and MCALens methods from a single tile of the simulated footprint. The KS and iKS maps have been smoothed with a Gaussian kernel with a standard deviation of $2$ arcmin, to render the structures more visible. Without smoothing, the KS and iKS maps are dominated by noise. The MCALens maps, however, are already noise-suppressed by the mass-mapping algorithm, so they do not require additional smoothing for visualisation purposes.
@@
The iKS map uses inpainting to fill in the missing data in the convergence field (see Sect. 3.1), effectively reducing the edge effects visible in the KS map. For reference, large masked regions are outlined in white in the plots of the reconstructed mass maps. Additionally, we see that the MCALens method removes noise, while keeping the stronger peaks in the convergence field, as expected from the method.
@@
where $i$ is the pixel index, $N_{\rm mask}$ is the total number of unmasked pixels (i.e., where $\mathbf{M}_{i}=1$), and $\mathbf{M}$ is the binary mask. The RMSE values reported in the table are computed as the mean over all independent realisations and all cosmologies. The table also lists the error on this mean RMSE. As shown in the table, applying Gaussian smoothing to the shear maps reduces the RMSE for the KS and iKS methods, indicating improved reconstruction accuracy. However, shear smoothing has minimal effect on the MCALens method, which consistently outperforms the others by achieving the lowest RMSE.
Table 1: RMSE comparison for mass map reconstruction
@@
KS (no $\gamma$ smoothing) $(1085\pm 2)\times 10^{-5}$
iKS (no $\gamma$ smoothing) $(1100\pm 2)\times 10^{-5}$
KS (with $\gamma$ smoothing) $(1018\pm 2)\times 10^{-5}$
iKS (with $\gamma$ smoothing) $(1023\pm 2)\times 10^{-5}$
MCALens (no $\gamma$ smoothing) $(978\pm 2)\times 10^{-5}$
MCALens (with $\gamma$ smoothing) $(976\pm 2)\times 10^{-5}$
222 The RMSE is evaluated between the true convergence map and the reconstructed mass maps obtained via the KS, iKS, and MCALens methods. It is evaluated for mass maps derived from shear maps that are either smoothed with a Gaussian kernel of standard deviation $0.4$ arcmin or left unsmoothed. The values represent the mean RMSE across all realizations of the 25 cosmologies along with the error on the mean. To ensure a fair comparison, each mass map is further smoothed with a Gaussian kernel optimized to minimize its respective RMSE.
Figure 1: Mass maps obtained using different reconstruction techniques, from a specific tile of the simulated footprint. Upper left: True convergence map derived directly from the simulated $\kappa$ field (after pixelising the per-object convergence). Upper right and lower left: Mass maps generated by the KS and iKS methods, respectively, both smoothed with a Gaussian kernel (standard deviation of 2 arcmin) for visualisation purposes. Lower right: MCALens mass map. The iKS method implements inpainting of the missing data, and the unmasked regions of the patch are outlined in white. Note that the white line outlines only the large unmasked areas, for better visualisation purposes.
@@ 4.3 Statistics calculation
Once the mass maps are generated, we calculated our HOS on $S/N$ maps, derived from the mass maps, as described in Section 3.2. We calculated the peak counts and wavelet peak counts only in regions where $\textbf{M}=1$, for both inpainted and non-inpainted maps, as the inpainted parts do not contain real signal. For both analyses, the $S/N$ maps were constructed using Eq. (3), and peaks were counted and binned into 20 linearly spaced bins over the range $-2<S/N<6$, following Ajani et al. (2020).
@@
For the mono-scale peak counts, we used a Gaussian kernel as the filter $\mathcal{W}$ in Eq. (3). We tested various kernel sizes and selected the size that yielded the best cosmological constraints for each method. The optimal kernel sizes were $\theta_{\rm ker}=2$ arcmin for the KS and iKS methods and $\theta_{\rm ker}=1$ arcmin for the MCALens method. We then counted the peaks in the resulting $S/N$ maps. Figure 2 displays the log-scaled histograms of the peak counts for the KS, iKS, and MCALens methods, averaged over all tiles and realisations for each of the 25 simulated cosmologies. The histograms are colour-coded as a function of the input $S_{8}$ showing the sensitivity of peak counts to cosmology as well as the difference between the mass-mapping methods. MCALens produces a significantly narrower peak distribution with a higher mode. This is a consequence of its de-noising process, which suppresses both noise and part of the signal, resulting in a lower-amplitude reconstructed $\kappa$ field compared to KS and iKS (see Fig. 1). However, the $S/N$ maps are computed using a noise level derived directly from the noise in the input shear field, Eq. (3), regardless of the mass-mapping method. As a result, the reduced amplitude of the MCALens $\kappa$ map translates into a peak distribution that is more concentrated around zero.
@@
Figure 2: Log-scaled histograms of the peak counts as a function of the signal-to-noise ratio. The three panels (from left to right) correspond to the KS, iKS, and MCALens methods, for the 25 simulated cosmologies. The histograms are colour-coded by their $S_{8}$ value.
@@
For the wavelet peak counts, we first decomposed the convergence maps into wavelet scales using the starlet transform, as described in Sect. 3.2. We used five wavelet scales, which correspond to resolutions $[2,4,8,16,32]$ arcmin, along with the coarse map. We then calculated the $S/N$ maps for each wavelet band separately, according to Eq. (3), and counted the peaks in each $S/N$ map. Figure 3 shows the log-scaled histograms of the wavelet peak counts for the KS, iKS, and MCALens methods, for the 25 simulated cosmologies. As expected, scales corresponding to smaller wavelet filters contain more peaks, as they capture smaller, more abundant structures in the convergence field. As in the mono-scale case, for most scales, MCALens results in a narrower distribution of peak counts, which again is a consequence of its de-noising approach that suppresses both noise and part of the signal.
@@
Figure 3: Log-scaled histograms of the wavelet peak counts as a function of the signal-to-noise ratio for the Kaiser-Squires (KS), iterative Kaiser-Squires (iKS), and MCALens methods, for the 25 simulated cosmologies. The four rows correspond to the three wavelet scales, $[8,16,32]$ arcmin and the coarse map, and the columns correspond to the three mass-mapping methods. The histograms are colour coded by their $S_{8}$ value.
@@ 4.4 Covariance matrices
We calculated the covariance matrices for the HOS, using the covariance training set described in Sect. 2.2. To avoid unphysical large-scale mode-coupling that arises from replicating the same light cone across tiles, we randomised the selection of light cones during the survey construction, following Harnois-Déraps et al. (2021). For each of the 124 full survey realisations, we drew the 19 tiles from different, randomly selected light cones. We applied this process to all 10 realisations and averaged them to create our final covariance matrix. The covariance matrix elements were calculated as
@@
where $N=1240$ is the number of realisations in the covariance training set, $x_{i}^{r}$ is the value of the statistic in the $i$-th bin of the $r$-th realisation, and $\bar{x}_{i}$ is the mean of the $i$-th bin over all realisations:
@@
To account for the finite number of bins and realisations, we used the de-biased estimator for an inverse Wishart distribution (Siskind 1972; Hartlap et al. 2007) for the inverse of the covariance matrix:
@@
where $n_{\rm bins}$ is the number of bins in the histogram of the HOS under consideration, and $\bm{\mathrm{C}}_{*}^{-1}$ is the inverse of the respective covariance matrix calculated according to Eq. (7). In Fig. 4, we show the correlation matrices for the single-scale peak counts, for the different mass-mapping methods. Fig. 5 shows the correlation matrices for the wavelet peak counts.
Figure 4: Correlation matrices for the single-scale peak counts, for the KS, iKS, and MCALens methods.
Figure 5: Correlation matrices for the wavelet peak counts, for the KS, iKS, and MCALens methods. The dashed white lines separate the different wavelet scales, which are positioned in increasing order of resolution from left to right, i.e. $[2^{\prime},4^{\prime},8^{\prime},16^{\prime},32^{\prime},\textrm{coarse}]$.